OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2

OpenStudy (anonymous):

4x + y = 8-------------- (A) x − 2y = 2 -------------- (B) 2*A+B 8x +2 y = 16 + x − 2y = 2 ------------ 9x=18 x=2 put x=2 in A 4*2+y=8 y=8-8 y=0

OpenStudy (anonymous):

ist u have to make the x or y value the same but with different signs (negative and positive). once u have done u cross out the same numbers, then u will be left with 2 like terms, u add them and also add the other side. simplify now. whatever is you value for x or y is, pick one of the equations above then substitute that value in to find the other one.

OpenStudy (anonymous):

i have the answer its 2 but i need the work

OpenStudy (anonymous):

ok. then what is your new equation since u made one of the x or y value the same

OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x − 3y = 3 5x − 4y = 3

OpenStudy (anonymous):

is your answer for x or y

OpenStudy (anonymous):

x

OpenStudy (anonymous):

pick one of the equation to find the y value. pick any

OpenStudy (anonymous):

4*(4x − 3y = 3) = = 16x-12y=12-----------A 3*(5x − 4y = 3) = = 15x-12y=9-----------B A-B 16x-12y=12 - 15x-12y=9 --------------- x=3

OpenStudy (anonymous):

thnx